OpenStudy (austinl):

Find an integrating factor and solve the given equation. \((3x^2y+2xy+y^3)+(x^2+y^2)y\prime=0\)

OpenStudy (anonymous):

Well in Order to use an integrating factor this differential equation must be in the form: \[\frac{ dy }{ dt } + P(t)y=g(t)\]

OpenStudy (anonymous):

Problem is I don't need to see how to isolate the y' into that form.

OpenStudy (austinl):

Nope, we are dealing with "exact" differential equations. Different sort of problem.... unfortunately.

OpenStudy (anonymous):

Nvm then. Sorry. Can't help.

OpenStudy (austinl):

Thanks anyway!

OpenStudy (austinl):

I have a couple ideas, but no-one is here to help.... I have a basic idea, but I am just so unsure of myself. *goes and cries in corner*

OpenStudy (accessdenied):

Sorry, I am not certain of Exact DE myself. Looking at some of my resources right now to remember how this one worked since seemed really weird in comparison to integrating factors / separable DE. :P

OpenStudy (austinl):

Ok... I have an idea. I believe the integrating factor is in the form, \(\dfrac{M_y(x,y)-N_x(x,y)}{n(x,y)}\mu\) \(M=3x^2y+2xy+y^3\) \(\dfrac{\partial M}{dy}(3x^2y+2xy+y^3)=3x^2+2x+3y^2\) \(N=x^2+y^2\) \(\dfrac{\partial N}{dx}(x^2+y^2)=2x\) So that means.... \(\dfrac{3x^2+2x+3y^2-2x}{x^2+y^2}\mu=\dfrac{3x^2+3y^2}{x^2+y^2}\mu\)

OpenStudy (accessdenied):

o.o Interesting. I didn't know that integrating factors existed for exact DE as well. That's something I never learned about. :) But I have a feeling with an answer like that you might be on the right track.

OpenStudy (austinl):

I made a small error... \(\dfrac{d\mu}{dx}=\dfrac{(3x^2+3y^2)\mu}{x^2+y^2}\) And I need to find \(\mu(x)\).

OpenStudy (austinl):

\(\dfrac{d\mu}{dx}=3\mu\) I have arrived here.....

OpenStudy (austinl):

Would that mean that \(\mu(x)=0\)?

OpenStudy (accessdenied):

that was unfortunate. got refreshed for some reason and lost reply. anyways is that specifically solving the DE or was there another factor involved. Since there are more solutions to that DE than mu = 0, generally exponential solutions, but I wasn't sure if there was an assumption to mu that made it mu = 0.

OpenStudy (austinl):

I am getting help from another site and they are saying that it would be an exponential of some sort, but I am unsure of how to even..... ugh.

OpenStudy (accessdenied):

Have you learned about separable DE?

OpenStudy (austinl):

I think so, but I am drawing a big blank right now.

OpenStudy (accessdenied):

basically, we would want to bring our mu's to one side and put the x's on the other \( \displaystyle \frac{d\mu}{dx} = 3 \mu \) \( \displaystyle \frac{1}{\mu} \frac{d\mu}{dx} = 3 \) (Assumes \(\mu \neq 0\) I've seen it written like \( \displaystyle \frac{1}{\mu} d \mu = 3 \; dx \) and integrated, although we could integrate both sides with respect to x and the dmu /dx x would sort of cancel out as well.

OpenStudy (accessdenied):

uh, not sure if the latex just stopped working for me only or its really not displaying properly at all.

OpenStudy (austinl):

I can see the \(\LaTeX\), but the site has been buggy for me today.

OpenStudy (accessdenied):

yea, same. anyways, does the process seem familiar what i wrote (in some invisible ink)?

OpenStudy (austinl):

Yeah, I remember now. I suppose I just need some caffeine of some kind.

OpenStudy (dumbcow):

hmm looks like this is not an Exact diff equ

OpenStudy (austinl):

It isn't, we are getting an integrating factor to make it an exact equation. Which has been calculated to be \(\mu(x)=e^{3x}\)

OpenStudy (dumbcow):

ahh i see just finished reading all the posts

OpenStudy (austinl):

Now I have, \(\mu(x)(3x^2y+2xy+y^3)+\mu(x)(x^2+y^2)y\prime=0\) Right? And that should make it exact.

OpenStudy (loser66):

yes.

OpenStudy (loser66):

I don't know why, but when you take \[\frac{M_y-N_x}{N}= 3\]you have the \(\mu = e^3x\)just plug it in and check, you have the problem exact, now, just solve as usual. I think so

OpenStudy (loser66):

*\(\mu = e^{3x}\)

OpenStudy (austinl):

\((3e^{3x}x^2y+2e^{3x}xy+e^{3x}y^3)+(e^{3x}x^2+e^{3x}y^2)y\prime=0\) ?

OpenStudy (austinl):

Holy crap this is gonna suck to solve.

OpenStudy (austinl):

\(C=(3x^2y+y^3)e^{3x}\) Does this look correct?

OpenStudy (dumbcow):

yes thats what i get

OpenStudy (austinl):

Very cool, that took much longer than I had hoped lol. Thanks @AccessDenied @Loser66 and @dumbcow Appreciate the help! ( ^_^)o自自o(^_^ )

OpenStudy (accessdenied):

(3x^2y + y^3)e^(3x) dp/dx = 6xy e^(3x) + 9x^2 y e^(3x) + 3y^3 e^(3x)= 3(2xy e^(3x) + 3x^2 y e^(3x) + y^3 e^(3x)) dp/dy = 3x^2 e^(3x) + 3y^2 e^(3x) = 3(x^2 e^(3x) + y^2 e^(x) ) dp/dx dx + dp/dy dy = 0 3(2xy e^(3x) + 3x^2 y e^(3x) + y^3 e^(3x)) + 3(x^2 e^(3x) + y^2 e^(3x)) dy/dx = 0 2xy e^(3x) + 3x^2 y e^(3x) + y^3 e^(3x) + (x^2 e^(3x) + y^2 e^(3x))dy/dx = 0 looks to be what we started with so yea, good job! Glad to be of some help! :)

OpenStudy (loser66):

Are you Chinese or Japanese or Korean?

OpenStudy (austinl):

Me? No... why?